# CrunchLLM: Multitask LLMs for Structured Business Reasoning and Outcome Prediction

### v1

@@ Abstract
Index Terms:
Large Language Models, Crunchbase dataset, Startup success prediction, Structured business reasoning, Multitask learning, Parameter-efficient fine-tuning, Prompt optimization, Explainable AI
@@ I Introduction
In brief, our main contributions are as follows:
  • We cast startup success prediction as a large language model (LLM) adaptation problem, leveraging Crunchbase data that combines structured company attributes with unstructured textual narratives.
  • We introduce CrunchLLM, a domain-adapted framework that integrates parameter-efficient fine-tuning and prompt optimization to specialize foundation models for entrepreneurship and innovation tasks.
  • We show that CrunchLLM achieves accuracy exceeding 80% in predicting startup success (IPO or acquisition), substantially outperforming both traditional machine learning methods and baseline LLMs.
  • Beyond predictive accuracy, CrunchLLM provides interpretable reasoning traces that justify predictions, thereby enhancing transparency and trust for use in venture investment, entrepreneurship research, and policy analysis.
@@ VI-A Models Evaluation
  • Qwen-3-8B - Mid-size 8B successor with improved reasoning/following; used as our strongest open-weight mid-tier reference for domain adaptation on Crunchbase.
@@ VI-B Metrics
We evaluate classification using Accuracy and F1, and justification generation using BERTScore (Precision, Recall, and F1). We perform a systematic evaluation of various models under three training regimes: zero-shot, few-shot, and full-data fine-tuning to assess their performance on multitask prediction and justification tasks.
The results in Table I confirm that full-data finetuning consistently outperforms zero and few-shot settings, both prediction and justification tasks. LLaMA-3-8B and LLaMA-3.2-3B follow closely, both achieving over 0.79 F1, and matching justification scores. Qwen-3-8B, when fully finetuned, achieves the best overall performance. Interestingly, GPT-2, despite being the smallest LLM evaluated, still outperforms XGBoost, demonstrating that language modeling pretraining transfers effectively to structured classification tasks, even with a modest parameter count. Note that XGBoost was only evaluated on the prediction task, as justification generation is inherently textual and cannot be performed by non-generative models. There is a strong correlation between a model’s prediction accuracy and its justification quality (BERT-F1). Models like Qwen-3-8B, which achieve high predictive F1, also generate semantically richer and more precise justifications, as indicated by its superior BERT-Precision. This supports the idea that accurate classification decisions guide better conditional text generation, which is vital in domains requiring interpretable AI.
A clear trend emerges when analyzing model size versus performance: smaller models, such as GPT-2 and TinyLLaMA, tend to plateau around 0.74–0.76 F1, even with full-data fine-tuning, while larger models in the 3B–8B parameter range consistently surpass 0.80 F1. This scaling advantage is also reflected in justification quality Qwen-3-8B, for instance, achieves over a 3% gain in BERT-F1 compared to GPT-2. These results emphasize that larger models not only offer better predictive power but are also more effective at multitask learning, producing both accurate classifications and coherent, contextually aligned justifications.
As seen in Table II, all models demonstrate substantial improvements in both F1 and Accuracy score when transitioning from the zero-shot setting to even 1k few-shot training examples. These improvements suggest that even minimal supervised exposure can dramatically bootstrap performance, especially for large pretrained models that benefit from in-context learning and minimal gradient updates. Moreover, the F1 score increases more rapidly than accuracy in early stages, which we attribute to class imbalance in the small training subsets. As previously discussed, minor distribution shifts in few-shot sampling can overrepresent certain classes, inflating F1 scores due to class-specific recall gains, even if overall accuracy ranking improves modestly.
TABLE I: Multitask Finetuning Model Performance Summary
Model Prediction Justification
F1 ACC BERT-Recall BERT-Precision BERT-F1
XGBoost 0.71 0.72 - - -
GPT-2 0.74 0.72 0.82 0.76 0.78
LLaMA-3-8B 0.81 0.82 0.82 0.76 0.78
LLaMA-3.2-3B 0.79 0.80 0.82 0.76 0.79
TinyLLaMA 0.76 0.78 0.81 0.75 0.78
Qwen-2.5-3B 0.74 0.77 0.81 0.79 0.80
Qwen-3-8B 0.90 0.89 0.80 0.83 0.81
TABLE II: Comparison between Zero-shot and Few-shot Performance
@@ VII-A Effect of Prompt Optimization
We systematically evaluated four prompt variants ranging from basic instructions (Prompt V1) to a fully optimized format (Prompt V4), comparing their impact on both prediction and justification performance. A clear trend emerges in Figure 4: as the prompt design improves, models yield higher performance across the board. The optimized prompt (Prompt V4) delivers the best F1 score of 0.87 and accuracy of 0.86 on the classification task, outperforming the initial prompt (Prompt V1) by over 5% absolute F1 improvement.
@@ Acknowledgements
This research is supported in part by the NSF under Grant IIS 2327113 and ITE 2433190, the NIH under Grants R21AG070909 and P30AG072946, and the National Artificial Intelligence Research Resource (NAIRR) Pilot NSF OAC 240219, and Jetstream2, Bridges2, and Neocortex Re- sources. We would like to thank NSF for support for the AI research resource with NAIRR240219. We thank the University of Kentucky Center for Computational Sciences and Information Technology Services Research Computing for their support and use of the Lipscomb Compute Cluster and associated research computing resources.

### v2

@@ I Introduction
In brief, our main contributions and novelties include the following:
  • We propose CrunchLLM, an adaptive LLM-based framework for predicting startup success from Crunchbase, integrating structured attributes with unstructured narratives.
  • We introduce prompt-guided multimodal fusion, a multi-task approach that dynamically combines structured tabular signals with company-profile text, enabling the model to identify prediction-relevant evidence from large-scale corpora, addressing challenges in existing LLMs that struggle with extracting key contents across modalities.
  • Our domain adaptation pipeline represents a novel strategy that integrates parameter-efficient fine-tuning and prompt optimization, tailored to entrepreneurship and innovation tasks where general-purpose LLMs typically underperform.
  • We demonstrate that CrunchLLM achieves state-of-the-art results on Crunchbase benchmarks, showing statistically significant gains over tabular-only, text-only, and generic LLM baselines (paired t-tests, $p<0.01$), while producing human-readable reasoning traces that enhance interpretability for investment, research, and policy-making.
@@ VI-B Metrics
We evaluate classification performance using Precision, Recall, F1, Accuracy (ACC), and Area Under the Curve (AUC). For justification generation, we adopt BERTScore (Precision, Recall, and F1) to assess the semantic quality of model explanations. We systematically benchmark multiple baselines across three training regimes: zero-shot, few-shot, and full-data fine-tuning.
The results in Table I and Table II confirm that full-data fine-tuning consistently yields the best performance across both prediction and justification tasks. Among baseline models, LLaMA-3-8B and LLaMA-3.2-3B achieve strong classification results (>>0.79 F1) with competitive justification quality. However, our proposed CrunchLLM substantially outperforms all baselines, surpassing F1 and Accuracy in prediction, while also delivering the highest AUC score. Notably, CrunchLLM also produces the most precise justifications, achieving superior BERT-Precision (0.83), reflecting its ability to ground explanations in meaningful business attributes.
Interestingly, GPT-2, despite being the smallest LLM evaluated, still surpasses XGBoost, showing that pretraining on language corpora transfers effectively to structured prediction. XGBoost, while competitive in classification (0.72 ACC), cannot generate textual justifications, underscoring the advantage of generative models in multitask reasoning. Overall, there is a clear correlation between a model’s predictive accuracy and the quality of its justifications: stronger classifiers also yield more coherent and semantically faithful explanations.
Table III further shows that all models benefit substantially from even minimal supervised exposure. Transitioning from zero-shot to 1k few-shot examples improves F1 and Accuracy by over 10% in most cases, with CrunchLLM demonstrating the largest gains. These improvements highlight the value of domain-specific supervision and our model’s ability to exploit it efficiently. We also observe that F1 increases more sharply than Accuracy at early stages, likely due to class imbalance in few-shot subsets that amplifies recall for minority classes.
Overall, the results show that scaling model size consistently improves both classification accuracy and justification quality, while prompt-guided multitask fine-tuning is essential for aligning predictions with coherent explanations. Our domain-adapted CrunchLLM establishes state-of-the-art performance by jointly leveraging structured company data and profile-style text, achieving both high predictive accuracy and interpretable reasoning. This dual capability demonstrates the novelty of our approach, as it provides a unified framework that advances beyond existing machine learning baselines and generic LLMs by delivering reliable predictions together with transparent, contextually grounded justifications for entrepreneurship, investment, and policy applications.
TABLE I: Prediction Performance Summary
Model Precision Recall F1 ACC AUC
XGBoost 0.70 0.72 0.71 0.72 0.75
GPT-2 0.73 0.75 0.74 0.72 0.78
LLaMA-3-8B 0.80 0.82 0.81 0.82 0.84
LLaMA-3.2-3B 0.78 0.80 0.79 0.80 0.84
TinyLLaMA 0.75 0.77 0.76 0.78 0.81
Qwen-2.5-3B 0.76 0.78 0.74 0.77 0.82
Our model 0.83 0.82 0.90 0.89 0.85
TABLE II: Justification Quality Summary
Model BERT-Recall BERT-Precision BERT-F1
GPT-2 0.82 0.76 0.78
LLaMA-3-8B 0.82 0.76 0.78
LLaMA-3.2-3B 0.82 0.76 0.79
TinyLLaMA 0.81 0.75 0.78
Qwen-2.5-3B 0.81 0.79 0.80
Our model 0.80 0.83 0.81
TABLE III: Comparison between Zero-shot and Few-shot Performance
@@ VII-A Effect of Prompt Optimization
We systematically evaluated four prompt variants ranging from basic instructions (Prompt V1) to a fully optimized format (Prompt V4), comparing their impact on both prediction and justification performance. A clear trend emerges in Figure 4: as the prompt design improves, models yield higher performance across the board. The optimized prompt (Prompt V4) delivers the best F1 score and accuracy on the classification task, outperforming the initial prompt (Prompt V1) by over 5% absolute F1 improvement.
@@ Acknowledgements
This research is supported in part by the NSF under Grant IIS 2327113 and ITE 2433190, the NIH under Grants R21AG070909 and P30AG072946, and the National Artificial Intelligence Research Resource (NAIRR) Pilot NSF OAC 240219, and Jetstream2, Bridges2, and Neocortex resources. We thank the University of Kentucky Center for Computational Sciences and Information Technology Services Research Computing for their support and use of the Lipscomb Compute Cluster and associated research computing resources.